Which of the following cells will produce the highest voltage? SHOW YOUR WORK.




a. Mn I Mn^2+ II Zn^2+ I Zn







b. Zn I Zn^2+ II Ni^2+ I Ni







c. Ni I Ni^2+ II Cu^2+ I Cu






Solve separately then write which one will produce the highest voltage.

1
Expert's answer
2022-06-13T17:02:10-0400



The overall cell potential can be calculated by using the equation E0cell=E0red−E0oxid.


a.

Mn - 2e- => Mn2+ E0 = -1,05 V

Zn2+ + 2e- => Zn E0 = -0,76 V



Mn + ZnCl2 => MnCl2 + Zn E0 = - 0,29


b.

Zn - 2e- => Zn2+ E0 = -0,76

Ni2+ + 2e- => Ni E0 = -0,23


NiCl2 + Zn => ZnCl2 + Ni E0 = -0,53


c.

Ni - 2e- => Ni2+ E0 = -0,23

Cu2+ -2e- = Cu E0 = 0,34


CuCl2 + Ni = NiCl2 + Cu E0 = -0,57


Answer: c.